Suppose the price of Fanta falls from R18 to R14 per bottle, and the quantity demanded rises from 250 to 400 bottles a day.  

 

1.1.1 Calculate the price elasticity of demand using the point elasticity formula.


1
Expert's answer
2021-11-07T19:40:28-0500

"PE=\frac{\frac{Q_2-Q_1}{Q_1}}{\frac{P_2-P_1}{P_1}}=\frac{\frac{400-250}{250}}{\frac{14-18}{18}}=-2.7."